From GlueXWiki
Jump to: navigation, search

November 17, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production update
    • Second package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at 126
  5. TDR (Lubomir)
  6. Other


Participants: Eugene, Bill, Dave, Chris, Simon, Vladimir, and Lubomir.


- Dave: second cell of the second package was installed. Third wire plane will be populated tomorrow by Chris and we will try to install the third cell by the end of the day. Working on the fourth and fifth cell cathodes. The fourth wire frame on which Dave replaced the caps, is being tested with HV by Anatoly (see below). It delayed the wire stringing by one week so far. The fifth wire frame was laminated.

- Discussions about the noise coming from a HV trace initiated by Eugene. We know the noise is coming from a HV trace on the back of the PCB. The amplitude of the noise is close to that of the standard signals, but is random with a frequency of <3kHz and width of ~20ns. So, the probability for interfering with the signals is <0.3% say for 1us window, so negligible. Starting with the sixth wire frame we are going to apply additional conformal coating (like the green coating on the PCBs) on the eight HV traces.

- Lubomir about the HV cap tests on the fourth wire frame: after replacing the caps last week and some conditioning we were able to get leakage current of 100-300nA initially but it later increased up to several uA. Anatoly replaced/fixed ~10 caps and the current went down to few tens of nA but still it was increasing after that. Normally, the initial current is high and then it goes down, opposite to what we see with this frame. This is the second frame on which we use 3kV rated caps and we had no issues like this with the first one. Also, Anatoly tested 37 of these 3kV caps separately and they didn't show such increase of the current with the time. Eugene: if the problem is in the 3kV caps we better wait for the 4kV capacitors (will have them hopefully in January!). We decided to continue with the HV tets: leave the frame under HV overnight and see if it will stabilize, then tomorrow morning will decide if we want to replace all the caps once more and then do more tests over the weekend. Then finally on Monday we should be in a position to decide if we will wait for the 4kV capacitors.

- Vladimir about the second cell tests: all channels working fine with a source. No noise on the second cell at the place where we found it on the other cells (as discussed above), but some kind of low frequency noise at a place close to another HV trace to be studied further.


- Bill: looking how to seal the cooling tube when using Fluorinert; will prototype with a many-fold. Rest of the gusset rings will be ready by January.

- Tests with DP460: samples were made with 2 mil and 0.5 mil Al-mylar glued on un-roughed g10. After one day curing, no evidence DP460 is better than Hysol we are using now. Will wait more and also do samples with roughed g10 and with Rohacell to repeat the CMU samples.

- Bill discussed also if we need to keep our equipment in 126 warm, since the heating will be down for a week during Christmas. We will not move the first package before Christmas. Beni should decide if some heating is needed for the equipment there.


- Chris finished testing card for the mass pre-amp tests. Made a cable to be used with the translator boards.

- LV cables: Anodes: spacing must be 2.75" (instead of 5") between the two connectors. Vladimir tested all the cathode groups with a mock-up cable. In two groups the 10-5-10" cable didn't work, but one was able to re-group the cards in such a way that 10-5-10" is long enough to be used everywhere. Prototype helped to understand that in most of the groups you have excessive length and only for a few you need the full length. Eugene insisted that we need to work more in optimizing the LV cabling by increasing the number of cable types so that we don;'t have excessive cables inside the detector. Bill and Vladimir will work on that.

- HV cabling: we need to start prototyping the cabling on the first package. Need the two connectors attached to the holding brackets and start soldering the individual wires to them. Chris needs schematics from Fernando.

Chamber testing at 126

- Eugene: when do you plan the testing to start? Main showstopper (apart form the fADC125) is the translator board. According to Chris it's not a problem to make 6 LV cables and two HV cables needed for the restricted (wires only) package tests.


- Slow progress, including now drawings from Bill, current version linked above. Lubomir will have a very preliminary version by Thanksgiving and Simon will continue working on it.


- Lubomir in vacations for three weeks after Thanksgiving. Will ask Beni for at least one meeting to be scheduled on Dec 8. If needed more will be scheduled.